OpenStudy (kaylaprincess):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. equation : f(x) = 3(x-1)^2+24

jimthompson5910 (jim_thompson5910):

f(x) = 3(x-1)^2+24 f(4) = 3(4-1)^2+24 ... replace every x with 4 f(4) = ???

jimthompson5910 (jim_thompson5910):

do the same for x = 7 to compute f(7) then you compute this expression to get the answer \[\Large \frac{f(7)-f(4)}{7-4}\]

OpenStudy (kaylaprincess):

alright thank you, question, when i square the binomial, do i just make it :(4-1)(4-1) or actually square it?

jimthompson5910 (jim_thompson5910):

I would do the stuff inside the parenthesis first, then square the result

jimthompson5910 (jim_thompson5910):

ex: (10-7)^2 = 3^2 = 9

jimthompson5910 (jim_thompson5910):

I'm following PEMDAS

OpenStudy (kaylaprincess):

oh yes of course, okay thank you

jimthompson5910 (jim_thompson5910):

no problem